OpenStudy (anonymous):

Write the equation of the quadratic function with roots 6 and 10 and a vertex at (8, 2).

OpenStudy (aum):

6 is a root implies (x-6) is a factor. 10 is a root implies (x-10) is a factor. Therefore the factors of the quadratic function are: (x-6)(x-10) The more general equation is: a * (x-6)(x-10) where 'a' is a constant. y = a * (x-6)(x-10) = a(x^2 - 16x + 60) (8,2) is a point on the parabola. Put x = 8 and y = 2 and solve for 'a'

OpenStudy (aum):

Another way to do this problem is: The equation of a parabola in the vertex form is: y = a(x-h)^2 + k where (h,k) is the vertex and 'a' is a constant. Here, (h,k) is (8,2). y = a(x - 8)^2 + 2 x = 6 is a root. That means when x = 6, y = 0. Substitute: 0 = a(6 - 8)^2 + 2 0 = 4a + 2 4a = -2 a = -1/2 y = -1/2 * (x - 8)^2 + 2 y = -1/2 * (x^2 - 16x + 64) + 2 y = -1/2x^2 + 8x - 32 + 2 y = -1/2 x^2 + 8x - 30 is the equation of the quadratic function.